Balls A and B accelerate from rest towards each other at 2 m/s2 and 3 m/s2 respectively. Given that the initial distance between the balls is 1500 m, determine:


i. the time taken for the balls to meet.

ii. the distance covered by ball A when the balls meet.

iii. the distance travelled by ball B when the balls meet.


1
Expert's answer
2021-06-14T14:49:55-0400

Gives

"a_1=2m\/sec"

"a_2=3m\/sec^2"

Distance (s)="1500m"

Ball A

Time take t1

"s=ut+\frac{1}{2}at^2"

s=1500-x

"u=0"

"t_1=\sqrt\frac{2s}{a_1}"

"t_1=\sqrt\frac{2(1500-x)}{2}"

Second ball

s=x

"t_2=\sqrt\frac{2\times x}{a_2}"

"t_1=t_2"

"\sqrt\frac{2(1500-x)}{2}=\sqrt\frac{2(x)}{3}"

Both side squre take

"1500-x=\frac{2x}{3}"

"4500-3x=2x"

"5x=4500"

"x=900m"

Both ball are meet same time

"t_1=t_2=t=\sqrt\frac{2\times900}{3}=\sqrt{600}sec"


Part (b)

Distance covered by ball A

"x'=1500-900=600m"

Part( c)

Dixtance covered ball B

"x=x''=900m"